### v2

# LOGIN: A Large Language Model Consulted Graph Neural Network Training Framework

Abstract.
Recent prevailing works on graph machine learning typically follow a similar methodology that involves designing advanced variants of graph neural networks (GNNs) to maintain the superior performance of GNNs on different graphs. In this paper, we aim to streamline the GNN design process and leverage the advantages of Large Language Models (LLMs) to improve the performance of GNNs on downstream tasks. We formulate a new paradigm, coined “LLMs-as-Consultants”, which integrates LLMs with GNNs in an interactive manner. A framework named L O G IN  (L LM c O nsulted G NN tra IN ing ) is instantiated, empowering the interactive utilization of LLMs within the GNN training process. First, we attentively craft concise prompts for spotted nodes, carrying comprehensive semantic and topological information, and serving as input to LLMs. Second, we refine GNNs by devising a complementary coping mechanism that utilizes the responses from LLMs, depending on their correctness. We empirically evaluate the effectiveness of L O G IN on node classification tasks across both homophilic and heterophilic graphs. The results illustrate that even basic GNN architectures, when employed within the proposed LLMs-as-Consultants paradigm, can achieve comparable performance to advanced GNNs with intricate designs. Our codes are available at https://github.com/QiaoYRan/LOGIN.
1. Introduction
Individual entities along with their respective interactions, typically denoting as $\langle{u},e_{u,v},{v}\rangle$, where $u$ and $v$ are entities and $e_{u,v}$ is the interaction between $u$ and $v$, can be effectively represented and analyzed in the form of graph-structured data. For instance, reciprocal following in social networks (Golbeck and Hendler, 2006; Mccallum et al., 2007), financial transfers between accounts (Akoglu et al., 2014; Pourhabibi et al., 2020), chemical bonds between molecules (Blundell, 1996; Kearnes et al., 2016), and citations of academic articles (Leskovec et al., 2005; Alcacer and Gittelman, 2006) can be generally represented as graph-like structures. It thus burgeons graph machine learning in various tasks for decades, such as recommendation (Ying et al., 2018; Ma et al., 2019; Basu et al., 1998), anomaly detection (Liu et al., 2021; Zhong et al., 2020; Krügel et al., 2002), drug discovery (Stärk et al., 2022; Wang et al., 2023b; Lo et al., 2018), and etc (Hamilton et al., 2017b; Riesen and Bunke, 2008; Broder et al., 2000). Recently, thanks to the advent of graph neural networks (GNNs) (Kipf and Welling, 2016; Hamilton et al., 2017a; Velickovic et al., 2017) and their demonstrated extensive capabilities, graph machine learning has become a highly active field attracting significant research attention.
Among these existing works, most of them strive to design different architectures of GNN to adapt to distinct graph types. The reason for that is the broad real-world sources of graph data induce considerable diversity across different graphs. One of the basic categorizations, for example, could be the case of homophily and heterophily (Rogers and Bhowmik, 1970a, b). Classic GNNs, due to their fundamental mechanisms of message passing and aggregation (Battaglia et al., 2018), have demonstrated superior performance on homophilic graphs but fail to generalize to heterophily scenarios where dissimilar nodes are connected (Zhu et al., 2020; Luan et al., 2020; Liu et al., 2022). Consequently, researchers have committed to developing various structures, aiming at improving the effectiveness of GNNs on particular types of graphs (Lim et al., 2021; Liu et al., 2022; Pei et al., 2019; Luan et al., 2020). In addition to individualized designs, there are also observed recent works applying Graph Neural Architecture Search (GNAS) techniques for seeking better GNN architecture from a series of essential components (Zhou et al., 2022; Zhang et al., 2023). Despite their various routes of the existing works, these methods hold a similar methodology: Design a specialized variant of graph neural networks to tackle specific analysis tasks.
However, designing these GNN variants to work effectively in real-world applications requires extensive manual experimentation. Researchers and practitioners may engage in a meticulous process of trial and error when adjusting model architectures. This process is usually time-consuming and demands a deep understanding of the theoretical underpinnings of GNNs and the practical requirements of business. Achieving a model that works and performs efficiently and accurately in specific scenarios often involves long periods of rigorous experimentation and iterative refinement. Moreover, these GNN variants tailored for different scenarios are not significantly different from the classic GNN models theoretically (Zhu et al., 2021). In other words, though they may incorporate modifications to address specific needs or optimize performance for particular graphs, the foundational principles and underlying architectures might remain largely consistent with classic GNNs. Whether the capacities of classic GNNs can be maximized for different scenarios has become an open research problem.
In this paper, our objective is to empower classic GNNs for consistently superior performance on graphs with varying characteristics. To this end, an influential off-shelf tool, the Large Lanuange Model (LLM), is adopted to help us achieve the goal. Following the scaling law (Kaplan et al., 2020), LLMs, such as 175B-parameter GPT-3 (Brown et al., 2020), display distinct behaviors compared to smaller pre-trained language models (PLMs) like the 330M-parameter BERT (Devlin et al., 2018) and the 1.5B-parameter GPT-2 (Radford et al., 2019). LLMs exhibit surprising emergent abilities (Wei et al., 2022), enabling them to solve complex tasks that smaller models struggle with (Zhao et al., 2023a). Consequently, the research community has focused increasingly on these LLMs, acknowledging their unique and advanced capabilities (Brown et al., 2020; Touvron et al., 2023; Alayrac et al., 2022; Liu et al., 2023; Ren et al., 2023; Mysore et al., 2023).
LLMs currently have not only achieved remarkable advancements in natural language processing (Brown et al., 2020; Touvron et al., 2023), but have also buoyed promising insights in other domains, such as computer vision (Alayrac et al., 2022; Liu et al., 2023, 2024; Reid et al., 2024) and recommendation systems (Ren et al., 2023; Mysore et al., 2023; Lin et al., 2024; Zhu et al., 2024), etc. This success can be attributed to the wealth of open-world knowledge stored in their large-scale parameters (Zhao et al., 2023a; Pan et al., 2024) and the emerging capabilities in logical reasoning (Huang and Chang, 2022; Brown et al., 2020). Analogously, we aim to leverage the advantages of LLMs to enhance graph machine learning, particularly the capacities of GNNs. We formulate a new paradigm that integrates LLMs with GNNs in an interactive manner, and we coin it “LLMs-as-Consultants”, following which, LLMs could explicitly contribute to the training process of GNN.
Under this paradigm, we propose a framework named L O G IN , short for L LM c O nsulted G NN tra IN ing , empowering the interactive utilization of LLMs within the GNN training process. As an interactive approach, the crucial issues of L O G IN lie in what GNNs should deliver to LLMs and how to feed the LLMs’ responses back to GNNs. First, for LLMs’ inputs, we delve into prompt engineering to craft concise prompts for spotted nodes, which carry comprehensive semantic and topological information. Second, to utilize the responses from LLMs, we devise a complementary coping mechanism depending on their correctness. Specifically, compared to ground truth labels, when LLMs predict correctly, we update node features to obtain semantic enhancement. Otherwise, we impute the misclassification to the potential presence of local topological noises, hence performing structure refinement. Besides, particular criteria may serve in the selection of essential nodes to reduce time complexity. In our implementation, we adopt GNN predictive uncertainty to assess the necessity of consulting LLMs.
To demonstrate the versatility and applicability of the proposed framework, we explore the effectiveness of L O G IN on node classification tasks across both homophilic and heterophilic graphs. By empirical studies, we illustrate that even classic GNNs, when employed within the proposed LLMs-as-Consultants paradigm, can achieve comparable performance to advanced GNNs with intricate designs.
The contributions of this paper can be summarized as follows.
  • To our knowledge, we are the first to propose the LLMs-as-Consultants paradigm of graph machine learning. Different from previous works, we integrate the power of LLMs interactively into GNN training.
  • Under this paradigm, we propose the L LM c O nsulted G NN tra IN ing  (L O G IN ) framework. This framework can be considered as a synthesis of previous methodologies, with a particularly tailored feedback strategy concerning the correctness of responses.
  • Experiments on six node classification tasks with both homophilic and heterophilic graphs demonstrate the effectiveness and generalizability of L O G IN .
The remainder of this paper is organized as follows. Section 2 surveys the related research in literature. Section 3 discusses the core differences and advantages of our proposed LLMs-as-Consultants paradigm, distinguishing it from previous methods. Section 4 introduces the concepts and preliminaries of this work. Section 5 details our L O G IN approach. Section 6 illustrates the evaluation of our model and baselines in six benchmark datasets, and Section 7 concludes the paper and discusses future work.
2. Related Works
2.1. GNN Variants
Recall that early proposed GNNs (Kipf and Welling, 2016; Hamilton et al., 2017a; Velickovic et al., 2017) are primarily designed for the most typical graph-structured data, such as citation networks, which are characterized as homophilic, homogeneous, and class-balanced graphs. Here homophily means nodes with similar attributes are more likely to connect (Hamilton, 2020; Rogers and Bhowmik, 1970a). For the second one, homogeneous graphs compose different types of entities (i.e., nodes) and relations (i.e., edges) (Wang et al., 2019; Zhang et al., 2019). Class balance refers to the situation where the distribution of classes or labels within a graph is nearly even (Zhao et al., 2021; Ju et al., 2024). Consequently, when generalized to heterophilic (Bo et al., 2021; Abu-El-Haija et al., 2019; Xu et al., 2018; Chen et al., 2020), heterogeneous (Zhang et al., 2019; Wang et al., 2019; Yun et al., 2019; Fu et al., 2020), and class-imbalanced graphs (Zhao et al., 2021; Qu et al., 2021; Park et al., 2021; Liu et al., 2021; Chen et al., 2021; Song et al., 2022), classic GNNs suffer from severe performance degradation.
However, the characteristics of heterophily, heterogeneity, and class imbalance are widely recognized in the graphs constructed from real-world scenarios. For example, heterophily is evident in online transaction networks as fraudsters are more likely to form connections with normal customers (Huang et al., 2022); in molecular graphs, different types of amino acids tend to link as to compose proteins (Zhu et al., 2020). Heterogeneous graphs can represent the various entities and their relations in recommendation (Shi et al., 2018; Hu et al., 2018) and cybersecurity systems (Hu et al., 2019; Hou et al., 2017). Class imbalance is present in fraud detection based on financial transaction graphs (Liu et al., 2021; Tian et al., 2023) due to the rarity of fraudulent cases.
To address the challenges GNNs face in generalizing to a wide range of real-world applications, researchers have been motivated to tailor classic GNNs to specific scenarios, namely heterophily, heterogeneity, class imbalance, and so on.
2.1.1. GNNs Designed for Heterophily
For heterophily, since similar nodes may not tend to bond, neighbor extension and GNN architecture refinement are two major designs for variants targeting heterophilic graphs. For neighbor extension, MixHop (Abu-El-Haija et al., 2019) and H2GCN (Zhu et al., 2020) enable GNNs to aggregate information across multi-hop ego-graphs. For GNN architecture refinement, H2GCN (Zhu et al., 2020) also excludes self-loop connections for non-mixing node embedding learning to produce distinguishable node embeddings. JKNet (Xu et al., 2018) combines intermediate representations from each layer, allowing flexible adaptation of neighborhood ranges to individual nodes. H2GCN (Zhu et al., 2020) and GCNII (Chen et al., 2020) further utilize all intermediate representations and only the first layer’s node embedding at each layer with the initial residual connection to realize inter-layer combination, respectively.
2.1.2. GNNs Designed for Heterogeneity
For heterogeneity, the key to designing a specified GNN variant is to propose a suitable aggregation function (Wang et al., 2022), which can better fuse the structural and semantic information of heterogeneous graphs. To name some, HAN (Wang et al., 2019) develops a hierarchical attention mechanism to determine the significance of various nodes and meta-paths, effectively capturing both the structural and semantic information within heterogeneous graphs. HetGNN (Zhang et al., 2019) employs bi-LSTM to aggregate neighbor embeddings, facilitating the learning of deep interactions among heterogeneous nodes. GTN (Yun et al., 2019) proposes an aggregation function that can automatically discover suitable metapaths during the message-passing process. MAGNN (Fu et al., 2020) designs three main components, i.e., node content transformation for input node attributes, intra-metapath aggregation for intermediate semantic nodes, and inter-metapath aggregation to integrate messages from multiple metapaths, with the aim to enhance its expressing power of heterogeneous graphs.
2.1.3. GNNs Designed for Class Imbalance
For class imbalance, the goal of designing GNN variants is to develop a GNN classifier that effectively handles both majority and minority classes (Ju et al., 2024). For example, GraphSMOTE (Zhao et al., 2021) utilizes synthetic minority oversampling in the embedding space to boost minority class representation. ImGAGN (Qu et al., 2021) and GraphENS (Park et al., 2021) address the class imbalance by generating synthetic minority nodes. PC-GNN (Liu et al., 2021) utilizes a label-balanced sampler to pick nodes for sub-graph training and a neighborhood sampler to balance the local topology. ReNode (Chen et al., 2021) and TAM (Song et al., 2022) adjust the weights of labeled nodes in the loss function utilizing topological information.
Theoretically, the aforementioned GNN variants tailored for different scenarios are not significantly different from the classic GNNs (Zhu et al., 2021). Practically, designing these variants to work is also challenging; it requires extensive manual experimentation over a long period. Our work aims to leverage LLMs to optimize GNNs, thereby enhancing the performance of GNNs on a diverse range of graphs, ultimately matching or even surpassing these intricately designed GNN variants.
2.2. LLMs for Graphs
The emergence of LLMs has inspired many explorations of utilizing LLMs for graph-structured data. Most of the works can be categorized into LLMs-as-Predictors and LLMs-as-Enhancers paradigms.
2.2.1. LLMs-as-Predictors
Representative methods following this paradigm, e.g. NLGraph (Wang et al., 2023a), GPT4Graph (Guo et al., 2023), GraphQA (Fatemi et al., 2023) and GraphText (Zhao et al., 2023b), attempt to harness the zero-shot and few-shot ability, along with the in-context learning ability to solve graph tasks by describing the graph structural topology in natural language. However, the complex global graph structure can hardly be compressed in a token-limited prompt, thereby utilizing LLMs solely only achieves the performance far from desired. Besides directly prompting LLMs, GraphLLM (Chai et al., 2023) conducts prefix tuning on an open-source LLM by concatenating graph-specific prefixes to the attention layers of a pre-trained LLM. InstructGLM (Ye et al., 2023) and GraphGPT (Tang et al., 2023) adapt LLMs for graph downstream tasks through instruction tuning, employing natural language and a graph-text aligner to express graph structural information, respectively. Nevertheless, neglecting the authenticated power of existing GNNs results in only moderate performance or substantial computing resource consumption.
2.2.2. LLMs-as-Enhancers
Different from the former paradigm, LLMs-as-Enhancers paradigm incorporates LLMs to enhance input graphs before GNN training. For example, TAPE (He et al., 2023) and Graph-LLM (Chen et al., 2023) prompt LLMs to interpret original texts attached to the nodes, and provide GNNs with semantically enhanced node features. SimTeG (Duan et al., 2023) parameter-efficiently finetunes (PEFT) LLMs with LoRA (Hu et al., 2021) on the textual corpus of graphs and then use the finetuned LLM to generate node representations for GNN prediction. LLM-TSE (Sun et al., 2023) explicitly instructs LLMs to produce a similarity score for two texts of two nodes, which subsequently leads to edge pruning. This methodology employs LLMs solely as a one-time data preprocessor before GNN training.
In summary, the LLMs-as-Predictors methods may result in suboptimal performance with higher resource consumption for overlooking the success of GNNs. The LLMs-as-Enhancers methods use LLMs to augment data before GNN training without integrating the LLM intelligence into GNN training. In comparison, our proposed LLM-as-Consultants paradigm aims to integrate LLMs into GNN training by consulting LLMs with a subset of nodes identified by GNNs as the most uncertain, in the attempt to enhance the classic GNNs with acceptable additional resource consumption. The core advancements of the LLMs-as-Consultants paradigm compared to previous paradigms, will be discussed detailedly in Section 3.
3. LLMs-as-Consultants Paradigm
We display the pipelines of three paradigms for integrating LLMs with graphs in Fig. 1. As shown in Fig. 1 (a) and (b), some pioneering research introduced in Section 2.2 has explored the potential of LLMs on graph data, which can be typically categorized into LLMs-as-Predictors and LLMs-as-Enhancers.
For the first kind, taking the node classification task as an example, some researchers harness the zero-shot and few-shot reasoning ability of LLMs to directly obtain node labels (Zhao et al., 2023b; Fatemi et al., 2023; Wang et al., 2023a; Guo et al., 2023). Others apply instruction tuning or prefix tuning to adapt LLMs specifically for graph tasks (Chai et al., 2023; Ye et al., 2023; Tang et al., 2023). However, most of them disregard the advantages of GNNs, and tuning LLMs may require substantial computing resources as well. For the latter one, LLMs-as-Enhancers, the LLMs are commonly utilized to enhance nodes’ semantic features (Chen et al., 2023; He et al., 2023; Duan et al., 2023) or to refine local topological structures (Sun et al., 2023). This kind of route is straightforward to apply LLMs in data preprocessing as a one-time enhancer, failing to couple the LLMs and GNNs interactively.
In contrast to existing approaches, our work introduces a novel paradigm that seamlessly integrates LLMs with GNNs through an interactive framework. We refer to this paradigm as ”LLMs-as-Consultants” (see Fig. 1 (c)). The fundamental distinction of our method lies in the recognition that LLMs can play an explicit and significant role in the training process of GNNs. This integration is not merely a superficial combination but an iteratively interactive manner where LLMs provide valuable insights and guidance that directly influence and enhance the learning capabilities of GNNs.
Figure 1. The paradigms for integrating LLMs with graphs.
4. Preliminaries
4.1. Text-Attributed-Graphs (TAGs)
Text-attributed graphs  (TAGs) are widely used in previous research on LLMs for graphs. A TAG can be formulated as:
(1) $\mathcal{G}=\left(\mathcal{V},\mathbf{A},\left\{s_{n}\right\}_{n\in\mathcal{V}}\right),$
where $\mathcal{V}$ is a set of $N$ nodes, $\mathbf{A}\in\{0,1\}^{N\times N}$ denotes the adjacency matrix, and $s_{n}\in\mathcal{D}^{L_{n}}$ is the text attached to node $n\in\mathcal{V}$, with $\mathcal{D}$ as the word dictionary, and $L_{n}$ as the text length. Note that our proposed framework is not limited to traditional TAGs that literally have texts as original attributes. In fact, most entities and their relations can be modeled and processed as graphs, and their characteristics can be expressed in text form.
4.2. Graph Neural Networks (GNNs)
For GNN training, the texts associated with the nodes should be encoded to an embedded space. We represent the node embeddings as $\mathbf{X}\in\mathbb{R}^{N\times D}$, in which each row $\boldsymbol{}{x}_{n}$ denotes the corresponding node embedding, with $D$ as its dimension number.
A general GNN framework includes information aggregation and update processes (Gilmer et al., 2017). The aggregation function is a differentiable function for collecting and combining information from a node’s neighbors to produce a summary vector. It typically involves a form of pooling or summing the features from neighboring nodes. This process can be defined generically as follows:
(2) $\mathbf{x}_{i}^{(k)}=\text{AGGREGATE}^{(k)}\left(\left\{\mathbf{x}_{j}^{(k-1)}:j\in\mathcal{N}(i)\right\}\right)$
Note that $x_{i}^{(k)}\in\mathbb{R}^{D}$ denotes the representation of node $i$ at layer $k$, with $\mathcal{N}(i)$ as the set of neighbors of node $i$. After aggregating the features from the neighbors, the update function combines the aggregated features with the node’s own features from the previous layer. The formula for updating is as follows:
(3) $\mathbf{x}_{i}^{(k)}=\text{UPDATE}^{(k)}\left(\mathbf{x}_{i}^{(k-1)},\mathbf{x}_{i}^{(k)}\right)$
Aggregation and update are fundamental operations in GNNs, enabling nodes to exchange information and update their representations based on the graph structure and node features. These operations facilitate the learning of complex patterns and dependencies in graph-structured data. For node classification, the final node representations from the last layer are used as input to a classifier for predicting the class labels.
4.3. Homophily Metrics
In this subsection, we introduce homophily metrics defined as follows to distinguish homophilic and heterophilic graphs.
Definition 4.1 (Graph-level Homophily Ratio).
Given a graph $\mathcal{G}=(\mathcal{V},\mathbf{A},\mathbf{X})$ and the node labels $\mathbf{Y}$, graph-level homophily ratio $h_{\mathcal{G}}$ is defined as the fraction of intra-class edges: $h_{\mathcal{G}}=\frac{|\mathcal{E}_{\mathrm{intra}}|}{|\mathbf{A}|}$, where $\mathcal{E}_{\mathrm{intra}}=\{(u,v)\mid A_{uv}=1\wedge\mathbf{y}_{u}=\mathbf{y}_{v}\}$.
Definition 4.2 (Node-level Homophily Ratio).
As to a node $v$ in $\mathcal{G}$, the node-level homophily ratio $h_{v}$ is defined as $h_{v}=\frac{|\mathcal{E}_{v}\cap\mathcal{E}_{\mathrm{intra}}|}{|\mathcal{E}_{v}|}$, where $\mathcal{E}_{v}=\{(u,v)\mid A_{uv}=1\}$ is the set of edges linked to $v$.
We utilize the aforementioned homophily metrics to categorize graph datasets and illustrate that our proposed L O G IN framework can help classic GNNs achieve remarkable performance on graphs exhibiting varying degrees of homophily.
5. Methodology
Figure 2. The pipeline of L
O
G
IN
5.1. Problem Formulation
In this paper, we explore the integration of LLM consultation into GNN training for the node classification task. We follow the transductive setting of node classification: Given some labeled nodes $\mathcal{V_{L}}\subset\mathcal{V}$ in graph $\mathcal{G}=(\mathcal{V},\mathbf{A},\left\{s_{n}\right\}_{n\in\mathcal{V}})$, we aim to classify the remaining unlabeled nodes $\mathcal{V_{U}}=\mathcal{V}\backslash\mathcal{V_{L}}$ within the same graph. Formally, the target is to learn a set of GNN parameters $\mathbf{W}$ to predict the unlabeled nodes with the guidance from LLMs:
(4) $f_{\mathbf{W}|\text{LLM}}:(\mathbf{A,\left\{s_{n}\right\}_{n\in\mathcal{V}}})\rightarrow\mathbf{Y}.$
Note that the ground truth labels of nodes, the pseudo labels predicted by GNNs, and the pseudo labels from LLMs are denoted as $\mathbf{Y},\mathbf{\hat{Y}},\mathbf{\hat{Y}}^{L}\in\{0,1\}^{N\times C}$ respectively, with $C$ as the number of classes.
5.2. Overview of L O G IN
Figure 2 demonstrates the framework of L O G IN . In this pipeline, we first identify target nodes that need consultation with LLM based on the GNN’s predictive uncertainty (c.f. Section 5.3). Then prompts consisting of the associated texts and local structures of these nodes are provided to LLMs for consultation (c.f. Section 5.4). After LLM consultation, a complementary mechanism is applied to fully utilize the LLMs’ responses, regardless of their classification correctness (c.f. Section 5.5).
5.3. Node Selection Based on Uncertainty
An intuitive node selection approach is to consult the LLM with all the nodes on the graph. However, such a manner fails to utilize the capacity of GNNs, which are proficient at handling simple nodes, and it could be inefficient due to the limited interaction speed with LLMs. Hence, we decide to consult LLMs only with partial difficult nodes in the graph, and adopt the variance of the predictions from several GNNs as a surrogate indicator for identifying the hard ones in our L O G IN .
Recall that Bayesian estimation is typically used to rate the predictive certainty in common neural networks (Kendall and Gal, 2017). Under the Bayesian framework, the conventionally fixed GNN parameters $\mathbf{W}$ are considered as random variables following specific distributions. The predictive probability of the Bayesian GNN with parameters $\mathbf{W_{b}}$ can be defined as Eq. (5).
(5) $p(\hat{\mathbf{Y}}\mid\mathbf{A},\mathbf{X})=\int_{\mathbf{W_{b}}}p(\hat{\mathbf{Y}}\mid\mathbf{W_{b}},\mathbf{A},\mathbf{X})p(\mathbf{W_{b}}\mid\mathbf{A},\mathbf{X})\mathrm{d}\mathbf{W_{b}}$
Since the true posterior $p(\mathbf{W}\mid\mathbf{A},\mathbf{X})$ in Eq. (5) is hard to calculate in practice, variational inference uses an arbitrary distribution $q_{\theta}(\mathbf{W})$ to approximate the posterior. Specifically, we use MC dropout variational inference, in which dropout could serve to perform variational inference. In this method, the variational distribution $q_{\theta}(\mathbf{W})$ is from a Bernoulli variable $\mathbf{M}_{\theta}$, representing whether the neurons are on or off, as Eq. (6) shows.
(6) $\mathbf{M}_{\theta}$ ${\sim\rm{Bernoulli}(\mathbf{\theta})},$
$q_{\theta}(\mathbf{W})$ $=p(\mathbf{W}\mid\theta)=p(\mathbf{W}_{\theta}),$
$\mathbf{W}_{\theta}$ $=\mathbf{M}_{\theta}\odot\mathbf{W}_{b}.$
where $\theta$ is the dropout rate, and $\mathbf{M}_{\theta}$ represents a binary mask that controls which neurons in GNNs are off. Then, $\{\hat{\mathbf{W}}_{t}\}_{t=1}^{T}$ is a collection of $T$ samples drawn from $\mathbf{W}_{\theta}$ in a Monte Carlo way. Through the minimization of the loss function defined in Eq. (7) using these weight samples, $\mathbf{W}_{b}$ can be acquired.
(7) $\mathcal{L}(\mathbf{W}_{b})=-\frac{1}{T}\sum_{t=1}^{T}\mathbf{Y}\log\left(f_{\hat{\mathbf{W}}_{t}}(\mathbf{A},\mathbf{X})\right)+\frac{1-\theta}{2T}\|\mathbf{W}_{b}\|^{2}$
After $\mathbf{W}_{b}$ is trained, the model uncertainty score $\mathbf{U}$, which is an $N$-dimension vector indicating the uncertainty score of each node, is calculated as in Eq. (8).
(8) $U(\hat{\mathbf{Y}}\mid\mathbf{A},\mathbf{X})$ $=\operatorname{Var}(\hat{\mathbf{Y}}\mid\mathbf{A},\mathbf{X})\approx\frac{1}{T}\sum_{t=1}^{T}\left(\hat{Y}_{t}-\frac{1}{T}\sum_{t=1}^{T}\hat{Y}_{t}\right)^{2}$
5.4. LLM Consultation
After we identify the nodes with the most uncertainty $\mathcal{V}_{uc}$ for LLM consultation, the next step in L O G IN is to figure out how to convert the semantic and topological information of these nodes into an understandable format by LLMs. We consult LLMs in a zero-shot node-by-node manner, i.e., constructing a single prompt for each uncertain node, without classification examples provided. Through this way, no fine-tuning method is involved, and the parameters of LLMs are fixed. Each node prompt is composed of three elements: instruction, input data, and output indicator. The Prompt Construction module in Fig. 2 shows a general prompt example, and we provide the detailed prompts in Appendix A.
5.4.1. Instruction.
The instruction explains the node classification task in the context of the graph type, depicting real-world meanings of nodes and edges, with a spectrum of the categories provided.
5.4.2. Input Data.
The input data includes information related to the target node $n$, namely its original text $s_{n}$, the two-hop neighborhood $\mathcal{N}_{2}(n)$ description, and the neighbor labels $\mathbf{Y_{\mathcal{N}_{2}(n)}},\mathbf{\hat{Y}_{\mathcal{N}_{2}(n)}}$ from both human annotations and GNN predictions. Instead of listing edges to express connectivity (Chen et al., 2023; Guo et al., 2023) , we opt to summarize one-hop and two-hop neighbors for the sake of semantic clarity and easier parsing. Note that we only select uncertain nodes from the train set, otherwise, it might bring a data leakage. The target node’s labels $y_{n},\hat{y}_{n}$ are also included in its corresponding prompt to provide more information for LLM predictions.
5.4.3. Output Indicator.
The output indicator is used to control the output format of LLMs for further analysis, specifically by offering a desired response example in JSON format. Regarding the response content, we aim for the output to include the most probable classification outcome $\hat{y}_{n}^{L}$ along with its explanation $e_{n}$.
5.5. LLM Response Feedback
After consulting LLMs with all selected uncertain nodes $\mathcal{V}_{uc}$, we have a group of LLM predicted pseudo labels and corresponding explanations $\left\{{n\in\mathcal{V}_{uc}}:(\hat{y}_{n}^{L},e_{n})\right\}$. Next, we aim to maximize the utilization of LLM responses and convert the information into signals suitable for GNN processing. The LLM responses can be divided into two types compared to ground truth labels, i.e., the correct predictions when $\hat{y}_{n}^{L}=y_{n}$ and the wrong predictions when $\hat{y}_{n}^{L}\neq y_{n}$. We not only utilize the correct predictions but also the misclassifications, and offer distinct approaches respectively. Incorporating such a complementary coping mechanism endows GNN training with full exploitation of LLM responses.
For clarity in expression, we denote the correct nodes and incorrect nodes as $\mathcal{V}_{r}$ and $\mathcal{V}_{w}$ respectively, satisfying $\mathcal{V}_{r}\cup\mathcal{V}_{w}=\mathcal{V}_{uc}$ and $\mathcal{V}_{r}\cap\mathcal{V}_{w}=\emptyset$.
5.5.1. When LLM is Right.
For a right node $n\in\mathcal{V}_{r}$, we append the explanation $e_{n}$ parsed from the LLM response to its original text $s_{n}$. And we attain its new node embedding $\mathbf{{x}_{n}}^{\prime}$ by encoding the new attached text from $e_{n}$:
(9) $\left\{\mathbf{{x}_{n}}^{\prime}\right\}_{n\in\mathcal{V}_{r}}\leftarrow\text{ENC}\left(\left\{s_{n}+e_{n}\right\}_{n\in\mathcal{V}_{r}})\right.$
5.5.2. When LLM is Wrong.
For those wrong nodes, the LLM responses are left out since the explanations for incorrect classifications could be unhelpful. Instead, we opt to attribute the misclassification to the topological information, e.g., complex structure patterns and potential adversarial noises. Therefore, we heuristically decide to simplify the local structure for the wrong nodes to make it easier for GNNs to learn. Specifically, around a wrong node $n\in\mathcal{V}_{w}$, we prune edges based on node similarity scores to denoise the local structure. As in representative graph structure learning methods, e.g. GNNGuard (Zhang and Zitnik, 2020), we measure similarity $d_{ni}$ between the features of the wrong node $n$ and its neighbor $i$ using cosine similarity:
(10) $d_{ni}=d^{cos}(\boldsymbol{x}_{n},~{}\boldsymbol{x}_{i})=\left(\boldsymbol{x}_{n}\odot\boldsymbol{x}_{i}\right)/\left(\left\|\boldsymbol{x}_{n}\right\|_{2}\left\|\boldsymbol{x}_{i}\right\|_{2}\right).$
Then, edges of node $n$ are pruned if their similarity scores fall below a user-defined threshold $d_{th}$. By pruning edges of the misclassified nodes $\mathcal{V}_{w}$ as in Eq. 11, we accomplished the topological refinement leveraging the wrong answers from LLMs.
(11) $\mathbf{A^{\prime}}\leftarrow\mathbf{1}{\left(\mathbf{D}-d_{th}\cdot\mathbf{I}\right)}\odot\mathbf{M}_{w}\odot\mathbf{A}\$
Here $\mathbf{D}=\left(d_{ij}\right)_{\mathbf{N}\times\mathbf{N}}$ denotes the similarity matrix, with $\mathbf{M}_{w}$ as a binary node mask representing $\mathcal{V}_{w}$, and $\mathbf{1}$ is a binary indicator.
In a nutshell, the response feedback stage is divided into two complementary scenarios. When the LLM classifies nodes correctly, we upgrade the original node embeddings with the encoded explanations. Conversely, we impute the misclassifications to the topological noise, hence pruning edges around the wrong nodes.
5.6. Train and Test
After the LLM response feedback stage, L O G IN is set to retrain GNN. When retraining, the GNN aggregates the original node features and the ones updated with the correct classification grounds. And for the nodes that LLMs misclassify, aggregation only happens along the retained edges after structure refinement. In this way, the message passing of GNNs is integrated with the LLM intelligence both semantically and topologically.
The GNN training process with one-time consultation within the L O G IN framework is illustrated in Algorithm 1. This process can be iteratively repeated until the maximum number of iterations is reached or the performance of the GNN has achieved an acceptable level.
During the test, the ground truth labels are inaccessible, so we only use the trained GNN without additional LLM consultation to make final predictions. Since we retrain the GNN over the graphs augmented by LLM consultation, i.e., with the node representations enhanced and the noisy links removed, the attained GNNs can perform aggregation in a more reliable neighborhood with more accurate semantic information and thus produce improvements.
Input: $\mathcal{G}=(\mathcal{V},\mathbf{A},,\left\{s_{n}\right\}_{n\in\mathcal{V}})$: A TAG, $\mathcal{V}_{\mathrm{train}}$: Set of training nodes, $\mathbf{X}$: Node representations, $\mathbf{Y}$: Labels of training nodes, $T$: Number of Monte Carlo samples, $\theta$: Dropout rate for uncertainty estimation, $\gamma$: Ratio of uncertain nodes, $d_{th}$: Similarity threshold
Output: The retrained GNN parameters after LLM consultation $\mathbf{W}_{r}$.
1
2Choose basic models for pretrained GNN $f_{\mathbf{W}_{p}}$ and retrained GNN $f_{\mathbf{W}_{r}}$;
3
4Initialize parameters $\mathbf{W}_{p}$ and $\mathbf{W}_{r}$;
5
6for $t=1,\ldots,T$ do
7       Sample $\hat{\mathbf{W}}_{t}$ from $\mathbf{W_{\theta}}$
8Train $\mathbf{W}_{p}$ and $\{\hat{\mathbf{W}}_{t}\}_{t=1}^{T}$ by minimizing Eq. (7);
9 Estimate the model uncertainty $U$ w.r.t. Eq. (8);
10 Divide $\mathcal{V}_{\mathrm{train}}$ into $\mathcal{V}_{\mathrm{c}}$ and $\mathcal{V}_{\mathrm{uc}}$ according to $U$ such that $\gamma=\frac{|\mathcal{V}_{\mathrm{uc}}|}{|\mathcal{V}_{\mathrm{train}}|}$;
11 for $n\in\mathcal{V}_{\mathrm{uc}}$ do
12       Consult the LLM with prompt $n$ and gain LLM-predicted label and its explanation $\hat{y_{n}}^{L},e_{n}$ ;
13
14      if $\hat{y_{n}}^{L}=y_{n}$ then
15             Update node features $\boldsymbol{x}_{n}$ w.r.t. Eq. (9);
16      else
17            Prune edges of $n$ w.r.t. Eq. (11);
18
19Retrain GNN $f_{\mathbf{W}_{r}}$ with new $\mathbf{A^{\prime}},\mathbf{X^{\prime}}$;
Return $\mathbf{W}_{r}$.
Algorithm 1 L
O
G
IN
6. Evaluation
In this section, we investigate the effectiveness of our L O G IN framework on both homophilic and heterophilic graph datasets, to address the following research questions:
  • RQ1: Does the L O G IN framework achieve performance comparable to state-of-the-art GNNs?
  • RQ2: How does the complementary coping mechanism for LLMs’ responses contribute to the L O G IN framework?
  • RQ3: How does L O G IN operate over specific nodes based on responses from LLMs?
  • RQ4: Can consulting more advanced LLMs in L O G IN unlock greater potential?
  • RQ5: How do models under the LLMs-as-Consultants paradigm perform compared with LLMs-as-Predictors and LLMs-as-Enhancers?
  • RQ6: Can consulting LLMs with more training nodes achieve higher performance increases with classic GNNs?
  • RQ7: Can L O G IN also help advanced GNNs enhance performance?
6.1. Experimental Setup
6.1.1. Datasets
We conducted extensive experiments on six datasets: three homophilic graphs and three heterophilic graphs, to demonstrate the versatility and applicability of our proposed L O G IN framework in handling graphs with distinct characteristics. For homophilic graphs, we collected the Cora (McCallum et al., 2000) and PubMed (Sen et al., 2008) datasets from widely used TAG benchmarks, while Arxiv-23 (He et al., 2023) was recently introduced to eliminate the data leakage unfairness when evaluating the impact of LLMs on graph learning.
For heterophilic graphs, we transformed the commonly-used web-page-link graphs: Wisconsin, Texas, and Cornell (Craven et al., 1998) into TAGs by sourcing and incorporating the raw texts [^1] http://www.cs.cmu.edu/~webkb/, which were not available previously in graph libraries. Besides, we fine-tuned DeBERTa-base (He et al., 2021) to encode raw texts into node embeddings. The basic statistics of the datasets are displayed in Table 1. The proportions of heterophilic nodes in graphs are calculated via Definition 4.2: If the node-level homophily score of node $n$ is below 0.5, then $n$ is defined as a heterophilic node.
Table 1. Statistics of the datasets.
Dataset #Node #Edge #Class #Feat %Heter
Cora 2708 5429 7 768 12%
PubMed 19717 44338 3 768 16%
Arxiv-23 46198 78548 40 768 8%
Wisconsin 265 938 5 768 79%
Texas 187 578 5 768 91%
Cornell 195 569 5 768 88%
6.1.2. Compared Methods
We integrate fundamental GNNs with L O G IN , namely GCN (Kipf and Welling, 2016), GraphSAGE (Hamilton et al., 2017a) and MixHop (Abu-El-Haija et al., 2019), to compare with the advanced state-of-the-art GNNs, including JK-Net (Xu et al., 2018), H2GCN (Zhu et al., 2020), APPNP (Gasteiger et al., 2018), GCNII (Chen et al., 2020), SGC (Wu et al., 2019) ,and SSP (Izadi et al., 2020), to demonstrate that the former can achieve performance on par with the latter. MLP, which predicts without adjacency information, also serves as a base learner combined with L O G IN to show the potential of our framework. We select Vicuna-v1.5-7b as our LLM consultant, an open-source LLM trained by fine-tuning Llama 2 on user-shared conversations.
For MLP, GCN, GraphSAGE, and MixHop, “OR” stands for the vanilla backbone model trained with original shallow node features, and “FT” denotes the vanilla backbone utilizing node embeddings encoded by the fine-tuned LM. “LO” represents the model trained within our L O G IN framework. Besides, in the ablation study, we refer to the feature update and structure refinement operations in the LLM response feedback stage as “F” and “S”, respectively.
In addition to advanced GNNs, we also compared L O G IN as an implementation of LLMs-as-Consultants paradigm, with the other existing paradigms, i.e. LLMs-as-Predictors and LLMs-as-Enhancers, respectively.
6.1.3. Implementations
We adopt accuracy on the test set of the nodes to evaluate the node prediction performance of all the listed models. We report the mean accuracy and standard error from five runs with varied random data splits, and all the reported results are statistically significant.
For hyper-parameters, in the uncertainty measure module, we implement Monte Carlo dropout variational inference by running models $T$ times with different neurons off, where $T$ represents the number of weight samples in MC dropout. $T$ is always set to 5 in our experiments. For the ratio $\gamma$ of uncertain nodes to consult, we adjust $\gamma$ slightly on each dataset around the proportion of heterophilic nodes shown in Table 1. In the response feedback stage, we tune the similarity threshold $s_{th}$ between $[0.1,0.2]$, according to the off-shell tool GNNGuard (Zhang and Zitnik, 2020).
6.2. Performance Comparison (RQ1)
Table 2. Performance comparison for node classification on homophilic and heterophilic graphs. The best results and the second best results among the classic GNNs with and without L
O
G
IN
and the advanced models are bold and underlined respectively.
Method Cora PubMed Arxiv-23 Wisconsin Texas Cornell
Classic GNNs MLP OR 0.6438 $\pm$ 0.0331 0.8805 $\pm$ 0.0032 0.6759 $\pm$ 0.0027 0.8113 $\pm$ 0.0718 0.8105 $\pm$ 0.0730 0.7538 $\pm$ 0.0669
FT 0.6897 $\pm$ 0.0102 0.9486 $\pm$ 0.0030 0.7789 $\pm$ 0.0023 0.8415 $\pm$ 0.0391 0.8211 $\pm$ 0.0681 0.8049 $\pm$ 0.0602
LO 0.7063 $\pm$ 0.0201 0.9505 $\pm$ 0.0036 0.7902 $\pm$ 0.0034 0.8528 $\pm$ 0.0588 0.8895 $\pm$ 0.0820 0.8051 $\pm$ 0.0618
GCN OR 0.8630 $\pm$ 0.0219 0.8635 $\pm$ 0.0083 0.6707 $\pm$ 0.0040 0.3736 $\pm$ 0.0672 0.4579 $\pm$ 0.0711 0.4308 $\pm$ 0.0664
FT 0.8683 $\pm$ 0.0191 0.9289 $\pm$ 0.0069 0.7624 $\pm$ 0.0051 0.4415 $\pm$ 0.1152 0.5526 $\pm$ 0.0832 0.5282 $\pm$ 0.0644
LO 0.8694 $\pm$ 0.0177 0.9396 $\pm$ 0.0030 0.7703 $\pm$ 0.0020 0.5057 $\pm$ 0.0430 0.5789 $\pm$ 0.0588 0.5231 $\pm$ 0.0292
GraphSAGE OR 0.8720 $\pm$ 0.0216 0.8849 $\pm$ 0.0026 0.6864 $\pm$ 0.0011 0.6113 $\pm$ 0.0662 0.5053 $\pm$ 0.0776 0.6051 $\pm$ 0.0389
FT 0.8592 $\pm$ 0.0363 0.9472 $\pm$ 0.0026 0.7881 $\pm$ 0.0019 0.7211 $\pm$ 0.1324 0.7579 $\pm$ 0.1123 0.7179 $\pm$ 0.1189
LO 0.8727 $\pm$ 0.0219 0.9511 $\pm$ 0.0036 0.7941 $\pm$ 0.0029 0.7434 $\pm$ 0.0930 0.7737 $\pm$ 0.1311 0.6872 $\pm$ 0.0896
MixHop OR 0.8601 $\pm$ 0.0281 0.8969 $\pm$ 0.0038 0.6774 $\pm$ 0.0029 0.5736 $\pm$ 0.1183 0.5526 $\pm$ 0.1500 0.4974 $\pm$ 0.0803
FT 0.8572 $\pm$ 0.0123 0.9493 $\pm$ 0.0030 0.7775 $\pm$ 0.0036 0.7092 $\pm$ 0.1035 0.7421 $\pm$ 0.1075 0.6718 $\pm$ 0.1397
LO 0.8624 $\pm$ 0.0253 0.9513 $\pm$ 0.0038 0.7818 $\pm$ 0.0040 0.7094 $\pm$ 0.0738 0.8158 $\pm$ 0.0930 0.7179 $\pm$ 0.0314
Advanced GNNs JK-Net 0.8579 $\pm$ 0.0001 0.8841 $\pm$ 0.0001 0.7532 $\pm$ 0.0012 0.7431 $\pm$ 0.0041 0.6649 $\pm$ 0.0046 0.6459 $\pm$ 0.0075
H2GCN 0.8692 $\pm$ 0.0002 0.8940 $\pm$ 0.0001 0.7382 $\pm$ 0.0011 0.8667 $\pm$ 0.0022 0.8486 $\pm$ 0.0044 0.8216 $\pm$ 0.0023
APPNP 0.8539 $\pm$ 0.0477 0.9355 $\pm$ 0.0060 0.7969 $\pm$ 0.0143 0.6830 $\pm$ 0.0470 0.7368 $\pm$ 0.0832 0.6410 $\pm$ 0.0480
GCNII 0.8833 $\pm$ 0.0027 0.7925 $\pm$ 0.0043 0.7847 $\pm$ 0.0068 0.7020 $\pm$ 0.0037 0.7135 $\pm$ 0.0039 0.7405 $\pm$ 0.0060
SGC 0.8509 $\pm$ 0.0648 0.8832 $\pm$ 0.0055 0.7740 $\pm$ 0.0160 0.5321 $\pm$ 0.0506 0.5526 $\pm$ 0.0811 0.4615 $\pm$ 0.0748
SSP 0.8616 $\pm$ 0.0289 0.9178 $\pm$ 0.0116 0.7976$\pm$ 0.0185 0.6302 $\pm$ 0.0850 $0$$\pm$ 0.0758 0.6923 $\pm$ 0.0314
To answer the first research question, we evaluate our proposed L O G IN framework on three homophilic datasets and three heterophilic datasets. The prediction accuracy scores and standard errors are reported in Table 2. We have the following observations.
6.2.1. Comparison with Vanilla Baselines.
Firstly, our method consistently outperforms the vanilla baselines trained with the original node features or the LM-finetuned node embeddings in most cases. There are only two exceptions in the Cornell dataset. Since Cornell is a relatively small dataset with only 39 nodes in the test set, as indicated by the standard errors, the experimental randomness is quite high with this small dataset. Nevertheless, L O G IN still helps improve the performance of MixHop on Cornell by 4.6% with a low deviation. Apart from this exception, all listed fundamental baselines trained within our L O G IN framework exceed the vanilla ones in node classification accuracy on the other five datasets, demonstrating the effectiveness of integrating LLMs as consultants into the GNN training process.
6.2.2. Comparison with Advanced GNNs.
Secondly, we are able to attain performance comparable to that of advanced GNNs by training fundamental models within the L O G IN framework. It is noteworthy that on PubMed and Texas, respectively known as benchmarks for homophilic and heterophilic graphs, we achieve the highest prediction accuracy among all the compared methods. This finding verifies the generalizability of our method on graphs with distinctive characteristics. For Cora, Arxiv-23, Wisconsin, and Cornell, L O G IN achieve remarkable performance on par with the SOTA GNNs such as H2GCN and SSP as well.
6.2.3. Comparison with the Simplest Model.
Thirdly, it draws our attention that regardless of the feature type and training paradigm we apply, MLP reveals great potential in the node classification task on heterophilic graphs. We believe that the TAGs with considerably high heterophily may be better regarded as natural language data rather than graph-structured data, for the links contained in these datasets seem to only carry noticeably limited information. Additionally, our L O G IN can clearly enhance MLP, inspiring further research on how to leverage simple existing tools to achieve superb capability with the help of the LLM-as-Consultants paradigm.
6.3. Ablation Study (RQ2)
To answer the second research question, we subdivide the response feedback stage into two distinct components, namely feature update and structure refinement. To show both of them contribute to our complementary coping mechanism for utilizing responses from LLMs, we remove these two modules respectively.
We present the results of ablation studies exclusively on the PubMed and Texas datasets, where L O G IN has demonstrated superior performance, as depicted in Fig. 3. Across both datasets, the complete L O G IN pipeline consistently achieves the highest performance, thus affirming the effectiveness and necessity of our complementary design for processing LLMs’ responses.
6.3.1. Feature Update.
Regarding the feature update component, we observe that it enhances prediction accuracy more effectively for a homophilic graph. Specifically, in the case of Cora, we notice from Fig. 3 (a) that L O G IN without feature updates results in a notable decrease in performance on certain occasions compared to the whole pipeline.
6.3.2. Structure Refinement.
Regarding the structure refinement component, L O G IN without edge pruning exhibits a significant performance decrease compared to the models with the complete coping mechanism, as shown in Fig. 3 (b). This observation aligns with our intuition, as in heterophilic graphs, the principal challenge for conventional GNNs in achieving generalization stems from the distinctiveness of their structural characteristics.
By removing each key component separately, we verify that regardless of the dataset type, in this case, the extent of heterophily, our design of the fault-tolerant complementary analysis strategy for LLMs’ responses is sound and necessary.
Figure 3. Ablation studies of L
O
G
IN
on PubMed and Texas.
6.4. Case Study (RQ3)
To answer the third research question, we select two individual nodes, respectively from Cora and Wisconsin, as specific examples to illustrate how L O G IN operates on them. These two nodes are both initially recognized as uncertain nodes and get misclassified by a pre-trained GNN. Through interaction with an LLM, the operation of feature enhancement or structure refinement is correspondingly conducted, thereby in turn helping the GNN make the right prediction.
6.4.1. From a Citation Graph: Cora.
We present node 356 from Cora as a representative example, whose ground truth label is Neural Networks. Unlike other papers, the title and abstract of this paper do not feature its label as a term explicitly, which also poses challenges for human classification. Besides, node 356 only has two one-hop neighbors, one of which is labeled differently as Probabilistic Method. This discrepancy may lead to failure in GNNs when processing this node. Nevertheless, thanks to the comprehensive understanding of the paper content facilitated by the parametric knowledge of LLMs, accurate prediction and a concise rationale are generated. Consequently, the subsequent semantic enhancement of node features significantly contributes to the final prediction of the GNN. The consultation process is elaborated in Fig. 4.
Figure 4. Case study of node 356 in Cora. Node 356 was first misclassified by GNN and selected as an uncertain node. The LLM consultation classified node 356 right along with a reasonable explanation. Then, the node representation was updated with the explanation. Finally, the retrained GNN managed to classify node 356 correctly.
6.4.2. From a Web-page-link Graph: Wisconsin.
Node 62 from Wisconsin represents a web page of a course, with content that is clear enough for humans to identify as a course homepage. However, due to its misleading neighborhood, where all nodes except itself in its 2-hop ego-graph do not represent course, among which 3 out of 4 are web pages of students, pre-trained GNNs cannot directly classify node 8 accurately. Moreover, the LLM consultant also provides the incorrect classification, along with an illogical rationale mentioning GNN prediction as in Fig. 5. This leads to edge pruning around node 62, which contributes to structure denoising that aids the re-trained GNN in making the right choice.
Figure 5. Case study of node 62 in Wisconsin. Node 62 was first misclassified by GNN and selected as an uncertain node. The LLM consultation classified node 62 wrong along with a self-contradictory explanation, which L
O
G
IN
left out. Then, edge pruning was performed around node 62. Finally, the retrained GNN managed to classify node 62 correctly.
The studies of two interesting cases we encountered in experiments highlight the solidity and efficacy of L O G IN when handling various scenarios, which are consistent with our motivation of designing an LLM-fault-utilized strategy for feedback.
6.5. Extensive Study (RQ4)
To answer the fourth question, we work on extended studies with more advanced LLMs, namely vicuna-v1.5-13b (Zheng et al., 2023) and GPT 3.5-turbo-0125 (Brown et al., 2020). Due to the constraints related to computational resources and OpenAI API calling, we provide results solely for Cora, based on two runs as presented in Table 3.
In this experiment, all fundamental baselines with GPT 3.5-turbo-0125 outperform the ones with open-source LLMs. Additionally, the implementation of L O G IN with vicuna-v1.5-13b demonstrates a modest enhancement in predictive accuracy compared with vicuna-v1.5-7b, which has fewer parameters. The trend in predictive accuracy aligns with our anticipations, indicating that the employment of a more advanced LLM within L O G IN framework indeed facilitates performance increase. This underscores the significant potential of the LLMs-as-Consultants paradigm when equipped with more powerful LLMs.
Table 3. Performance comparison among L
O
G
IN
with different LLMs on Cora.
LLMs MLP GCN GraphSAGE
Vicuna-v1.5-7b 0.7063 $\pm$ 0.0331 0.8694 $\pm$ 0.0102 0.8727 $\pm$ 0.0201
Vicuna-v1.5-13b 0.7202 $\pm$ 0.0201 0.8702 $\pm$ 0.0191 0.8739 $\pm$ 0.0135
GPT 3.5-turbo-0125 0.8123 $\pm$ 0.0254 0.8992 $\pm$ 0.0099 0.8856 $\pm$ 0.0102
6.6. Comparison among LLM-based Paradigms (RQ5)
To answer the fifth research question, we investigate the LLMs-as-Predictors, LLMs-as-Enhancers, and our LLM-as-Consultants paradigms by testing typical methods derived from each. For LLMs-as-Predictors, we prompt vicuna-v1.5-7b (Zheng et al., 2023) to directly get predictions. For LLMs-as-Enhancers, we adopt the TAPE method (He et al., 2023) equipped with llama2-13b-chat(Touvron et al., 2023).
Table 4. Performance comparison among LLMs-as-Predictors, LLMs-as-Enhancers and LLMs-as-Consultants paradigms.
Dataset Method LLMs-as-Predictors LLMs-as-Enhancers LLMs-as-Consultants
(vicuna-v1.5-7b) (TAPE + llama2-13b-chat) (L
O
G
IN
+ vicuna-v1.5-7b)
Cora MLP 0.7432 $\pm$ 0.0131 0.7675 $\pm$ 0.0187 0.7343 $\pm$ 0.0841
GCN 0.8630 $\pm$ 0.0101 0.8759 $\pm$ 0.0151
GraphSAGE 0.8625 $\pm$ 0.0093 0.8699 $\pm$ 0.0167
PubMed MLP 0.7847 $\pm$ 0.0632 0.9475 $\pm$ 0.0046 0.9508 $\pm$ 0.0040
GCN 0.9257 $\pm$ 0.0063 0.9401 $\pm$ 0.0043
GraphSAGE 0.9464 $\pm$ 0.0033 0.9505 $\pm$ 0.0031
Arxiv-23 MLP 0.7547 $\pm$ 0.0927 0.7905 $\pm$ 0.0041 0.7909 $\pm$ 0.0064
GCN 0.7751 $\pm$ 0.0029 0.7734 $\pm$ 0.0028
GraphSAGE 0.7935 $\pm$ 0.0029 0.7961 $\pm$ 0.0029
Note that the accuracy scores in Table 4 differ from those in Table 2, since we take the results of TAPE + llama2-13b-chat concerning their paper, which are reported for four runs. In Table 4, we also display the results of the same four data splits [^2] https://github.com/XiaoxinHe/TAPE to guarantee the comparison is fair.
As Table 4 shows, the LLMs-as-Consultants paradigm consistently outperforms the LLMs-as-Predictors paradigm across all datasets. Additionally, compared to the LLMs-as-Enhancers paradigm, our method surpasses the TAPE method equipped with llama2-13b-chat with only one exception, despite it employs an open-source LLM with significantly more parameters and prompts it with all nodes rather than a small subset selected. Our LLMs-as-Consultants paradigm demonstrates greater compatibility with lower time and resource consumption.
6.7. Consulting with More Nodes (RQ6)
To address the sixth research question, we conducted experiments on the Cora and Wisconsin datasets using MLP and GraphSAGE to investigate the impact of consulting LLMs with more nodes. Our findings demonstrate that incorporating more nodes in the consultation process with LLMs can lead to significant performance improvements. The results are in Table 5.
Through these experiments, we increased the number of nodes involved in the LLM consultation by 10% each. The results on two example homophilic and heterophilic graphs show that as the number of consulted nodes increases, there are corresponding improvements in the overall performance of the model. This suggests that LLMs, when provided with a broader context through more extensive node consultation, can offer more insightful responses, thereby refining the model’s predictions and decisions. These findings underscore the potential benefits of leveraging LLMs more comprehensively, highlighting the effectiveness of the node consultation process to fully exploit the capabilities of LLMs in improving GNN performance.
Table 5. Performance comparison of consulting with more nodes on Cora and Wisconsin.
Method Cora Wisconsin
MLP+LOGIN 0.7063 ± 0.0331 0.8528 ± 0.0588
MLP+LOGIN+10% nodes 0.7147 ± 0.0318 0.8605 ± 0.0493
GrraphSAGE+LOGIN 0.8727 ± 0.0219 0.7434 ± 0.0930
GraphSAGE+LOGIN+10% nodes 0.8789 ± 0.0310 0.7527 ± 0.0698
6.8. L O G IN with Advanced GNNs (RQ7)
To address the seventh research question, we performed experiments incorporating APPNP and GCNII with L O G IN to demonstrate that L O G IN can also enhance the performance of advanced GNNs.
The results in Table 6 indicate notable performance gains on both Cora and Wisconsin. These enhancements underscore the versatility and broad applicability of L O G IN in enhancing advanced GNNs as well, reinforcing its potential as a general solution for boosting various types of GNNs. These experiments provide evidence that L O G IN is generally effective and not limited to improving specific types of GNNs.
Table 6. Performance comparison of APPNP and GCNII trained with or without the L
O
G
IN
framework.
Method Cora Wisconsin
APPNP 0.8539 ± 0.0477 0.7969 ± 0.0143
APPNP+LOGIN 0.8614 ± 0.0332 0.7973 ± 0.0202
GCNII 0.8833 ± 0.0027 0.7847 ± 0.0068
GCNII+LOGIN 0.8879 ± 0.0105 0.7903 ± 0.0121
7. Conclusion and Future Work
In this work, we propose a new paradigm of leveraging LLMs for graph tasks, coined “LLMs-as-Consultants”. Following this paradigm, our L O G IN framework empowers interactive LLM consultation in the GNN training process. We identify uncertain nodes in the GNN pre-training stage, prompt LLMs with rich semantic and topological information compression, and parse LLMs’ responses in a not only fault-tolerant but also fault-utilized way to enhance GNN re-training. Extensive experiments on both homophilic and heterophilic graphs illustrate the validity and versatility of our proposed LLMs-as-Consultants paradigm.
For future endeavors, we intend to further refine GNN weights during the training phase using insights from LLMs’ responses, building upon our current research. Additionally, how to efficiently involve a large scale of nodes in LLM consultation to improve the scalability of the LLMs-as-Consultants paradigm remains an open question.
Moreover, a significant challenge that remains is determining how to efficiently involve large-scale nodes in the LLM consultation process. This is crucial for improving the scalability and practicality of the LLMs-as-Consultants paradigm. Addressing this issue involves developing methods that can handle the extensive computational requirements and ensure that the consultation process remains effective and responsive even as the number of nodes increases. Exploring solutions such as distributed computing, optimized query strategies, and hierarchical consultation frameworks could be potential pathways to making this paradigm scalable and efficient for real-world applications.
[…]
Appendix A Example Prompts
We display 2 example node prompts on datasets Cora and Wisconsin respectively in Table 7 for reproducibility.
Table 7. Example Prompts for Cora and Wisconsin.
Datasets Prompts
Cora Given a citation graph: { “id”:“cora_356”, “graph”: { “node_idx”: 356, “node_list”: [356, 190, 510, 519, 2223, 192], “one_hop_neighbors”: [190, 510], “two_hops_neighbors”: [192, 519, 2223], “node_label”: [“Neural_Networks”, “Neural_Networks”, “Probabilistic_Methods”, “Neural_Networks”, “Neural_Networks”, “Neural_Networks”] }}, where the 0th node is the target paper, with the following information: Title: [Title Text], Abstract: [Abstract Text]. And in the ‘node label’ list / ‘GNN-predicted node label’ list, you’ll find the human-annotated / GNN-predicted subcategories corresponding to the neighbors within two hops of the target paper as per the ‘node_list’.
Question: Which CS sub-category does this paper belong to? Give the most likely CS sub-categories of this paper directly, choosing from ”Case_Based”, ”Genetic_Algorithms”, ”Neural_Networks”, ”Probabilistic_Methods”, ”Reinforcement_Learning”, ”Rule_Learning”, ”Theory”. Ensure that your response can be parsed by Python json, using the following format as an example: {”classification result”: ”Genetic_Algorithms”, ”explanation”: ”your explanation for your classification here”}. Ensure that the classification result must match one of the given choices.
Wisconsin Given a webpage link graph: { “id”: “wisconsin_62”, “graph”: { “node_idx”: 62, “node_list”: [62, 166, 189, 165, 84], “one_hop_neighbors”: [166, 189], “two_hops_neighbors”: [84, 165], “node_label”: [“course”, “student”, “student”, “student”, “faculty”] }}, where the 0th node is the target webpage, with the following content: [webpage content text]. And in the ‘node label’ list / ‘GNN-predicted node label’ list, you’ll find the human-annotated / GNN-predicted subcategories corresponding to the neighbors within two hops of the target paper as per the ‘node_list’.
Question: Which category does this webpage belong to? Give the most likely category of this webpage directly, choosing from “course”, “faculty”, “student”, “project”, “staff”. Ensure that your response can be parsed by Python json, using the following format as an example: {“classification result”: “student”, “explanation”: “your explanation for your classification here”,}. Ensure that the classification result must match one of the given choices.
Appendix B more baselines of different LLMs-for-Graphs Paradigms
To further show the effectiveness of the LLMs-as-Consultants paradigm, we conducted more experiments on Arxiv (Lim et al., 2021) dataset with more baselines of both LLMs-as-Predictors and LLMs-as-Enhancers paradigms as Table 8 shows. Note that Arxiv is a semi-homophilic graph with 169,343 nodes and 31,166,24 edges, among which 37% nodes are heterophilic.
GraphGPT (Tang et al., 2023) and InstructGLM (Ye et al., 2023) are LLMs-as-Predictors methods. The results of GraphGPT and InstructGLM on Arxiv dataset are directly borrowed from their papers.
SimTeG (Duan et al., 2023) and TAPE (He et al., 2023) are both LLMs-as-Enhancers methods. SimTeG involved LLM fine-tuning and TAPE prompted LLM for predictions and explanations to enhance the node features. For SimTeG, we trained GNNs with the released embeddings of Arxiv. For TAPE, we reproduced the prompting with Vicuna-7b-v1.5 to make the comparison fair. Note that SimTeG finetuned LLMs on all the training nodes and TAPE enhanced every node embedding, while our method only consults LLMs with some uncertain nodes, requiring less time and memory, but still performs on par with the baselines.
Table 8. Performance comparison with more baselines of LLMs-as-Predictors and LLMs-as-Enhancers paradigm on Arxiv.
Paradigm Method Acc
LLMs-as-Predictors GraphGPT 0.7511
LLMs-as-Predictors InstructGLM 0.7570 ± 0.0012
LLMs-as-Enhancers MLP+SimTeG 0.7432 ± 0.0023
LLMs-as-Enhancers MLP+TAPE 0.7461 ± 0.0029
LLMs-as-Consultants MLP+LOGIN 0.7476 ± 0.0036
LLMs-as-Enhancers GraphSAGE+SimTeG 0.7624 ± 0.0050
LLMs-as-Enhancers GraphSAGE+TAPE 0.7558± 0.0029
LLMs-as-Consultants GraphSAGE+LOGIN 0.7609± 0.0041